# Computation method for perfectoid purity and perfectoid BCM-regularity

### v4

@@ Abstract.
In this paper, we introduce a criterion for a sufficient condition of perfectoid purity and a computation method for perfectoid pure threshold of ${\rm div}(p)$.
@@ 1. Introduction
Recently, the theory of singularities in mixed characteristic has been developed in many papers [MSTWW], [BMPSTWW24], [HLS]. The theory has applications for both the theory of algebraic geometry and the ring theory in mixed characteristic, for example, the minimal model theory [BMPSTWW23], [TY23], and the Skoda-type theorem [HLS]. In this paper, we study the class of singularities perfectoid purity [p-pure] and perfectoid BCM-regularity [bcm-reg], which are known as analogues of singularities in positive characteristic $F$-purity and strongly $F$-regular, respectively. It is known that for a Noetherian local ring $(R,\mathfrak{m})$ with $p\in\mathfrak{m}$, if $R$ is complete intersection and $R/pR$ is $F$-pure or strongly $F$-regular, then $R$ is perfectoid pure or perfectoid BCM-regular, respectively. On the other hand, we have a useful criterion called “Fedder’s criterion” [Fedder], [Glassbrenner] for $F$-purity and strongly $F$-regularity of complete intersection local rings, thus many examples can be constructed combining them.
In this paper, we construct new examples of perfectoid pure or perfectoid BCM-regular rings. Let $(R,\mathfrak{m})$ be a $p$-torsion free Noetherian local ring such that $p\in\mathfrak{m}$ and $\overline{R}:=R/pR$ is $F$-finite. We say that $R$ is perfectoid pure if there exists a pure extension $R\to B$ to a perfectoid. If $R$ is perfectoid pure, the perfectoid pure threshold $\operatorname{ppt}(R;{\rm div}(p))$ of ${\rm div}(p)$ is defined by
$\inf\left\{i/p^{e}\geq 0\middle|\begin{array}[]{l}\textup{there exist a perfectoid $R$-algebra $B$ and $\varpi$ a $p^{e}$-power}\\ \textup{root of $p$ such that $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure}\end{array}\right\},$
see Proposition 2.9 for details. It is known [p-pure]*Theorem 6.6 that if $R$ is complete intersection and $\overline{R}$ is $F$-pure, then $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1$ (cf. Remark 2.10). The first purpose of the paper is to construct examples of perfectoid pure rings $R$ with $\operatorname{ppt}(R;{\rm div}(p))\neq 1$. For this purpose, we define the ring homomorphism $\Phi_{R,n}\colon\overline{R}\to\overline{W}_{n}(R):=W_{n}(R)/pW_{n}(R)$, where $W_{n}(R)$ is the ring of Witt vectors over $R$ of length $n$, by
$\Phi_{R,n}\colon\overline{R}\to\overline{W}_{n}(R)\ ;\ a\mapsto(\widetilde{a}^{p},0,\ldots,0),$
where $\widetilde{a}\in R$ is a lift of $a$ in $R$ (see Proposition 4.1 for well-definedness). We regard $\overline{W}_{n}(R)$ as an $\overline{R}$-module via $\Phi_{R,n}$ and denote it by $Q_{R,n}$, then we note that $Q_{R,1}=F_{*}\overline{R}$. We introduce the following notions.
@@
  • Let $n$ be a positive integer. We say that $R$ is $n$-quasi-$F$-split if $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$ splits as $\overline{R}$-modules and $R$ is quasi-$F$-split if $R$ is $m$-quasi-$F$-split for some positive integer $m$.
  • The quasi-$F$-splitting height $\mathrm{ht}(R)$ is defined by
@@
  • Let $n,e$ be positive integers. We define $\overline{R}$-module $Q_{R,n,e}$ by the following pushout diagram as $\overline{R}$-modules;
@@
    and the induced $\overline{R}$-module homomorphism is denoted by
    $\Phi_{R,n,e}\colon R\xrightarrow{\Phi_{R,n}}Q_{R,n}\to Q_{R,n,e}.$
  • We say that $R$ is quasi-$(F,F^{\infty})$-split if $R$ is quasi-$F$-split and $\Phi_{R,\mathrm{ht}(R),e}$ splits as $\overline{R}$-modules for every positive integer $e\geq 1$.
The first main result in this paper is Theorem A.
@@
We assume $R$ is complete intersection. Let $n$ be a positive integer.
@@
    (Theorem 5.15). In particular, if $R/pR$ is quasi-$F$-split, then $R$ is perfectoid pure (Theorem 5.12).
@@
    $R$ is quasi-$(F,F^{\infty})$-split of height $n$ if and only if $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1-1/p-\cdots-1/p^{n-1}$ (Corollary 5.16).
Furthermore, we obtain a Fedder type criterion for quasi-$F$-splitting (Theorem 4.12) and obtain examples as follows (cf. Example 5.9).
@@
The next main result Theorem B is on the perfectoid BCM-regularity ([bcm-reg]*Definition 6.9) for graded rings.
@@
Moreover, we obtain a Fedder type criterion for quasi-$F$-splitting (Theorem 4.12), thus Theorem 4.12 is a criterion for perfectoid BCM-regularity for rings satisfying the assumption in Theorem B in characteristic two. By the inversion of adjunction type result (Proposition 4.6), if $\overline{S}$ is quasi-$F$-split, so is $S$. Furthermore, we construct examples as follows.
Theorem C.
@@
    We set $\mathfrak{m}:=(S_{>0},p)$ then $S_{\mathfrak{m}}$ is perfectoid BCM-regular.
@@
Finally, we compute a perfectoid pure threshold for quasi-$F$-split graded rings with $a(\overline{S})=0$.
Theorem D.
@@
Furthermore, we obtain an interesting example of non-perfectoid pure ring as follows.
@@
A key ingredient of all main results is running a program (Definition 5.4) to obtain a sequence of elements of $H^{d}_{\mathfrak{m}}(\overline{R})$. Whether or not this program stops is closely related to the singularity of a given ring (cf. Remark 5.5(3) and Proposition 5.7).
@@ Acknowledgements.
The author wishes to express his gratitude to Ryo Ishizuka, Teppei Takamastu, Kenta Sato, and Tatsuro Kawakami for valuable discussion. He is also grateful to Linquan Ma for helpful comments. The author was supported by JSPS KAKENHI Grant number JP20J11886.
@@ 2.1. Notation
In this subsection, we summarize the notation and terminologies used in this paper.
@@
    Throughout this paper, we fix a prime number $p$ and we set $\mathbb{F}_{p}:=\mathbb{Z}/p\mathbb{Z}$.
@@
    For a ring $R$ of characteristic $p>0$, $F:R\to R$ denotes the absolute Frobenius ring homomorphism, that is, $F(r)=r^{p}$ for every $r\in R$.
@@
    For a ring $R$ of characteristic $p>0$, we set $R_{perf}:=\mathrm{colim}_{F}R$ and call it by the perfection of $R$.
@@
    Throughout this paper, we use the terminology perfectoid by means of [BMS19]*Definition 3.5. For a semiperfectoid $R$, the perfectoidazation is denoted by $R_{perfd}$ (cf. [BS]*Section 7).
@@
    Let $R$ be a ring and $f$ an element of $R$. A sequence $\{f^{1/p^{e}}\}_{e\geq 0}$ is called a compatible system of $p$-power roots of $f$ if $(f^{1/p^{e+1}})^{p}=f^{1/p^{e}}$ for every $e\geq 0$ and $f^{1/p^{0}}=f$. Furthermore, the ideal $(f^{1/p^{e}}\mid e\geq 0)$ of $R$ is denoted by $(f^{1/p^{\infty}})$.
@@
    For a ring homomorphism $f\colon A\to B$ and a finitely generated ideal $I$ of $A$, we say that $f$ is $I$-completely faithfully flat if for every $I$-torsion $A$-module $N$, we have $\operatorname{Tor}_{i}^{A}(B,N)=0$ for every $i\neq 0$ and $B/IB$ is faithfully flat over $A/I$. If $I=(p)$, we call it by $p$-completely faithfully flat.
@@ 2.2. The ring of Witt vectors
We take a ring homomorphism $\pi\colon A:=\mathbb{Z}[X_{f}\mid f\in R]\to R$ such that $\pi(X_{f})=f$, then $\pi$ is surjective. Since $\pi$ induces a ring homomorphism $W(\pi)\colon W(A)\to W(R)$ and $A$ is $p$-torsion free, we may assume that $R$ is $p$-torsion free. We take $\varphi\colon W(R)\to\prod_{\mathbb{Z}_{\geq 0}}R$ as in Definition 2.2, then $\varphi$ is injective. Let $s\colon\prod_{\mathbb{Z}_{\geq 0}}R\to\prod_{\mathbb{Z}_{\geq 0}}R$ be a map defined by $s(a_{0},a_{1},\ldots)=(0,a_{0},a_{1},\ldots)$, then we have $s(\alpha\beta)=s(\alpha)s(\beta)$ for $\alpha,\beta\in\prod_{\mathbb{Z}_{\geq 0}}R$. Let us show (1). We have
@@
for $a,b\in R$. Since $\varphi$ is injective, we obtain the assertion (1). Next, let us show (2). Furthermore, we have $\varphi\circ V(\alpha)=ps\circ\varphi(\alpha)$ for $\alpha\in W(R)$, thus
@@ 2.3. Test perfectoid and perfectoid pure threshold
In this subsection, we introduce the notion of test perfectoid and perfectoid pure threshold of ${\rm div}(p)$. The existence of the test perfectoids is proved in Appendix A.
@@
Let us show (1). We take a perfectoid $R$-algebra $B$, then replacing $B$ by a $p$-completely faithfully flat extension of $B$, we may assume $B$ is $R_{\infty}$-algebra. By [CS]*Proposition 2.1.11(b), the $p$-adic completion $V^{\prime}:=B\widehat{\otimes}_{R_{\infty}}R^{\prime}_{\infty}$ of $B\otimes_{R_{\infty}}R^{\prime}_{\infty}$ is a perfectoid. Since $R_{\infty}\to R^{\prime}_{\infty}$ is $p$-completely faithfully flat, so is $B\to B^{\prime}$ by Remark 2.5. Since there exists an $R$-algebra homomorphism $R^{\prime}_{\infty}\to B^{\prime}$, as desired. Next, let us show (2). We take a perfectoid $R/I$-algebra $B$. Since $R_{\infty}$ is a test perfectoid, there exists a $p$-completely faithfully flat extension $B\to B^{\prime}$ of perfectoid $R$-algebras such that there exists an $R$-algebra homomorphism $R_{\infty}\to B^{\prime}$. Since $B^{\prime}$ is an $R/I$-algebra and $R_{\infty}$-algebra, it is an $R_{\infty}/IR_{\infty}$-algebra. Thus, we obtain the $R/I$-algebra homomorphism $(R_{\infty}/IR_{\infty})_{perfd}\to B^{\prime}$, as desired. ∎
@@
Next, let us show assertion (2). By definition, we have
@@ 3. Witt vectors of rings with Frobenius lifts
(cf. [Ill79]*Section 1.3) We use Notation 3.1. There exists a ring homomorphism
@@
We use Notation 3.1. We regard $W_{n}(A)$ as $A$-module by $s_{\phi}$ for every positive integer $n$. Then $V\colon\phi_{*}W_{n-1}(A)\to W_{n}(A)$ is an $A$-module homomorphism for $n\geq 2$.
@@
We use Notation 3.1. We define a map
@@
We use Notation 3.1. Maps $\Delta_{W_{n}}$ are group homomorphisms and satisfy the following properties
@@
(c.f. [kty]*Lemma 3.17) We use Notation 3.1. Let $a\in A$. We define the map
@@
=$\overset{(\star_{1}}{=}\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)-\sum^{n-1}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)$
@@
We use Notation 3.1. We regard $W_{n}(A)$ as $A$-module by $s_{\phi}$. Then we obtain an $A$-module isomorphism
@@
We use Notation 3.1. The $A$-module homomorphism $s_{\phi}\circ\phi\colon A\to\phi_{*}W_{n}(A)$ induces $\overline{s_{\phi}\circ\phi}\colon A/pA\to\phi_{*}W_{n}(A)/pW_{n}(A)$. Then $\overline{s_{\phi}\circ\phi}(\overline{a})=[a^{p}]$ for $a\in A$ and $\overline{a}$ is the image of $a$ by $A\to A/pA$.

### v5

@@ Abstract.
In this paper, we introduce the notion of quasi-$F$-splitting for rings in mixed characteristic. By comparing quasi-$F$-splitting with perfectoid purity, we obtain a new inversion of adjunction-type result. Furthermore, we study the possible values of the perfectoid pure threshold of ${\rm div}(p)$ and construct new examples of perfectoid pure rings.
@@ 1. Introduction
The theory of singularities in mixed characteristic has recently seen significant development in numerous papers, such as [MSTWW], [BMPSTWW24], and [HLS]. This theory has found applications in both algebraic geometry and commutative algebra in mixed characteristic, including the minimal model program [BMPSTWW23], [TY23], and Skoda-type theorems [HLS].
In this paper, we study two classes of singularities: perfectoid purity [p-pure] and perfectoid BCM-regularity [bcm-reg], which are regarded as analogues of $F$-purity and strong $F$-regularity in positive characteristic, respectively. It is known that for a Noetherian local ring $(R,\mathfrak{m})$ with $p\in\mathfrak{m}$, if $R$ is a complete intersection and $R/pR$ is $F$-pure or strongly $F$-regular, then $R$ is perfectoid pure or perfectoid BCM-regular, respectively. On the other hand, a useful criterion known as Fedder’s criterion [Fedder], [Glassbrenner] provides a practical method to test $F$-purity and strong $F$-regularity for complete intersection local rings. Many examples have been constructed using this criterion.
We now fix notation and recall the definition of perfectoid purity. Let $(R,\mathfrak{m})$ be a $p$-torsion free Noetherian local ring with $p\in\mathfrak{m}$ and such that $\overline{R}:=R/pR$ is $F$-finite. We say that $R$ is perfectoid pure if there exists a pure extension $R\to B$ to a perfectoid ring. If $R$ is perfectoid pure, the perfectoid pure threshold $\operatorname{ppt}(R;{\rm div}(p))$ of ${\rm div}(p)$ is defined as
$\inf\left\{i/p^{e}\geq 0\middle|\begin{array}[]{l}\text{there exist a perfectoid $R$-algebra $B$ a $p^{e}$-th root $\varpi$}\\ \text{of $p$, such that the map $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure}\end{array}\right\},$
see Proposition 2.9 for details. It is known [p-pure]*Theorem 6.6 that if $R$ is a complete intersection and $\overline{R}$ is $F$-pure, then $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1$ (cf. Remark 2.10).
The first goal of this paper is to obtain a new inversion of adjunction-type result. To this end, we define the ring homomorphism
$\Phi_{R,n}\colon\overline{R}\to\overline{W}_{n}(R):=W_{n}(R)/pW_{n}(R),\quad a\mapsto(\widetilde{a}^{p},0,\ldots,0),$
where $W_{n}(R)$ is the ring of Witt vectors over $R$ of length $n$, and $\widetilde{a}\in R$ is a lift of $a$ (see Proposition 4.1 for well-definedness). We view $\overline{W}_{n}(R)$ as an $\overline{R}$-module via $\Phi_{R,n}$ and denote it by $Q_{R,n}$. Note that $Q_{R,1}=F_{*}\overline{R}$.
We introduce the following notions:
@@
  • Let $n$ be a positive integer. We say that $R$ is $n$-quasi-$F$-split if $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$ splits as $\overline{R}$-modules. We say that $R$ is quasi-$F$-split if it is $n$-quasi-$F$-split for some $n$.
  • The quasi-$F$-splitting height $\operatorname{ht}(R)$ is defined by
@@
    and we set $\operatorname{ht}(R):=\infty$ if $R$ is not quasi-$F$-split.
  • For positive integers $n$ and $e$, define the $\overline{R}$-module $Q_{R,n,e}$ as the pushout in the following diagram:
@@
    The induced $\overline{R}$-module homomorphism is denoted by
    $\Phi_{R,n,e}\colon\overline{R}\xrightarrow{\Phi_{R,n}}Q_{R,n}\to Q_{R,n,e}.$
  • We say that $R$ is quasi-$(F,F^{\infty})$-split if $R$ is quasi-$F$-split and $\Phi_{R,\operatorname{ht}(R),e}$ splits as $\overline{R}$-modules for every $e\geq 1$.
Our first main result establishes a connection between quasi-$F$-splitting and perfectoid purity:
@@
Assume $R$ is a complete intersection and let $n\geq 1$.
@@
    In particular, if $R/pR$ is quasi-$F$-split, then $R$ is perfectoid pure.
@@
    $R$ is quasi-$(F,F^{\infty})$-split of height $n$ if and only if $R$ is perfectoid pure and
    $\operatorname{ppt}(R;{\rm div}(p))=1-\frac{1}{p}-\cdots-\frac{1}{p^{n-1}}.$
We also establish a Fedder-type criterion for quasi-$F$-splitting (Theorem 4.12) and provide explicit examples, including cases where $\operatorname{ppt}(R;{\rm div}(p))\neq 1$ (cf. Example 5.13).
@@
In addition, we prove a result on the possible values of the perfectoid pure threshold:
@@
(Corollary 5.15) We assume $R$ is complete intersection and perfectoid pure. If
$\operatorname{ppt}(R;{\rm div}(p))>\frac{p-2}{p-1},$
then $\operatorname{ppt}(R;{\rm div}(p))$ satisfies
for some positive integer $n$.
The next main result Theorem C concerns the perfectoid BCM-regularity ([bcm-reg]*Definition 6.9) for graded rings.
Theorem C.
@@
Moreover, we establish a Fedder-type criterion for quasi-$F$-splitting (Theorem 4.12), which in particular serves as a criterion for perfectoid BCM-regularity for rings satisfying the assumptions of Theorem C in characteristic two. Furthermore, an inversion of adjunction-type result (Proposition 4.6) shows that if $\overline{S}$ is quasi-$F$-split, then so is $S$.
We now present applications of the above theorems by constructing explicit examples.
Theorem D.
@@
    We set $\mathfrak{m}:=(S_{>0},p)$. Then $S_{\mathfrak{m}}$ is perfectoid BCM-regular.
@@
We also compute the perfectoid pure threshold for quasi-$F$-split graded rings with $a(\overline{S})=0$.
Theorem E.
@@
Finally, we present an example of a ring that is perfectoid pure, but whose self-tensor product is not.
@@ Acknowledgements.
The author wishes to express his gratitude to Ryo Ishizuka, Teppei Takamastu, Kenta Sato, and Tatsuro Kawakami for valuable discussion. He is also grateful to Linquan Ma for helpful comments. The author was supported by JSPS KAKENHI Grant number JP24K16889.
@@ 2.1. Notation
In this subsection, we summarize the notation and terminology used throughout this paper.
@@
    We fix a prime number $p$ and set $\mathbb{F}_{p}:=\mathbb{Z}/p\mathbb{Z}$.
@@
    For a ring $R$ of characteristic $p>0$, we denote by $F\colon R\to R$ the absolute Frobenius ring homomorphism, defined by $F(r)=r^{p}$ for all $r\in R$.
@@
    For a ring $R$ of characteristic $p>0$, we define the perfection of $R$ as $R_{\mathrm{perf}}:=\mathrm{colim}_{F}R$.
@@
    We use the terminology perfectoid as defined in [BMS19]*Definition 3.5. For a semiperfectoid ring $R$, its perfectoidization is denoted by $R_{\mathrm{perfd}}$ (cf. [BS]*Section 7).
@@
    Let $R$ be a ring and $f$ an element of $R$. A sequence $\{f^{1/p^{e}}\}_{e\geq 0}$ is called a compatible system of $p$-power roots of $f$ if $(f^{1/p^{e+1}})^{p}=f^{1/p^{e}}$ for all $e\geq 0$ and $f^{1/p^{0}}=f$. Furthermore, the ideal $(f^{1/p^{e}}\mid e\geq 0)$ of $R$ is denoted by $(f^{1/p^{\infty}})$.
@@
    For a ring homomorphism $f\colon A\to B$ and a finitely generated ideal $I$ of $A$, we say that $f$ is $I$-completely faithfully flat if for every $I$-torsion $A$-module $N$, we have $\operatorname{Tor}_{i}^{A}(B,N)=0$ for every $i\neq 0$ and $B/IB$ is faithfully flat over $A/I$. If $I=(p)$, we say that $f$ is $p$-completely faithfully flat.
@@ 2.2. The ring of Witt vectors
We take a ring homomorphism $\pi\colon A:=\mathbb{Z}[X_{f}\mid f\in R]\to R$ such that $\pi(X_{f})=f$, then $\pi$ is surjective. Since $\pi$ induces a ring homomorphism $W(\pi)\colon W(A)\to W(R)$ and $A$ is $p$-torsion free, we may assume that $R$ is $p$-torsion free. We take $\varphi\colon W(R)\to\prod_{\mathbb{Z}_{\geq 0}}R$ as in Definition 2.2, then $\varphi$ is injective. Let $s\colon\prod_{\mathbb{Z}_{\geq 0}}R\to\prod_{\mathbb{Z}_{\geq 0}}R$ be a map defined by $s(a_{0},a_{1},\ldots)=(0,a_{0},a_{1},\ldots)$, then we have $s(\alpha\beta)=s(\alpha)s(\beta)$ for $\alpha,\beta\in\prod_{\mathbb{Z}_{\geq 0}}R$. We prove assertion (1). We have
@@
for $a,b\in R$. Since $\varphi$ is injective, we obtain assertion (1). Next, we prove assertion (2). Furthermore, we have $\varphi\circ V(\alpha)=ps\circ\varphi(\alpha)$ for $\alpha\in W(R)$, thus
@@ 2.3. Test perfectoid and perfectoid pure threshold
In this subsection, we introduce the notion of test perfectoid and perfectoid pure threshold. The existence of the test perfectoids is proved in Appendix A.
@@
We prove (1). We take a perfectoid $R$-algebra $B$, then replacing $B$ by a $p$-completely faithfully flat extension of $B$, we may assume that $B$ is $R_{\infty}$-algebra. By [CS]*Proposition 2.1.11(b), the $p$-adic completion $V^{\prime}:=B\widehat{\otimes}_{R_{\infty}}R^{\prime}_{\infty}$ of $B\otimes_{R_{\infty}}R^{\prime}_{\infty}$ is a perfectoid. Since $R_{\infty}\to R^{\prime}_{\infty}$ is $p$-completely faithfully flat, so is $B\to B^{\prime}$ by Remark 2.5. Since there exists an $R$-algebra homomorphism $R^{\prime}_{\infty}\to B^{\prime}$, as desired. Next, we prove (2). We take a perfectoid $R/I$-algebra $B$. Since $R_{\infty}$ is a test perfectoid, there exists a $p$-completely faithfully flat extension $B\to B^{\prime}$ of perfectoid $R$-algebras such that there exists an $R$-algebra homomorphism $R_{\infty}\to B^{\prime}$. Since $B^{\prime}$ is an $R/I$-algebra and $R_{\infty}$-algebra, it is an $R_{\infty}/IR_{\infty}$-algebra. Thus, we obtain the $R/I$-algebra homomorphism $(R_{\infty}/IR_{\infty})_{perfd}\to B^{\prime}$, as desired. ∎
@@
Next, we prove assertion (2). By definition, we have
@@ 3. Witt vectors of rings with Frobenius lifts
(cf. [Ill79]*Section 1.3) We use the notation introduced in Notation 3.1. There exists a ring homomorphism
@@
We use the notation introduced in Notation 3.1. We regard $W_{n}(A)$ as $A$-module by $s_{\phi}$ for every positive integer $n$. Then $V\colon\phi_{*}W_{n-1}(A)\to W_{n}(A)$ is an $A$-module homomorphism for $n\geq 2$.
@@
We use the notation introduced in Notation 3.1. We define a map
@@
We use the notation introduced in Notation 3.1. Maps $\Delta_{W_{n}}$ are group homomorphisms and satisfy the following properties
@@
(c.f. [kty]*Lemma 3.17) We use the notation introduced in Notation 3.1. Let $a\in A$. We define the map
@@
=$\overset{(\star_{1})}{=}\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)-\sum^{n-1}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)$
@@
We use the notation introduced in Notation 3.1. We regard $W_{n}(A)$ as $A$-module by $s_{\phi}$. Then we obtain an $A$-module isomorphism
@@
We use the notation introduced in Notation 3.1. The $A$-module homomorphism $s_{\phi}\circ\phi\colon A\to\phi_{*}W_{n}(A)$ induces $\overline{s_{\phi}\circ\phi}\colon A/pA\to\phi_{*}W_{n}(A)/pW_{n}(A)$. Then $\overline{s_{\phi}\circ\phi}(\overline{a})=[a^{p}]$ for $a\in A$ and $\overline{a}$ is the image of $a$ by $A\to A/pA$.